# ParallelComp: Parallel Long-Context Compressor for Length Extrapolation

### v1

@@ Abstract
Efficiently handling long contexts is crucial for large language models (LLMs). While rotary position embeddings (RoPEs) enhance length generalization, effective length extrapolation remains challenging and often requires costly fine-tuning. In contrast, recent training-free approaches suffer from the attention sink phenomenon, leading to severe performance degradation. In this paper, we introduce ParallelComp, a novel training-free method for long-context extrapolation that extends LLMs’ context length from 4K to 128K while maintaining high throughput and preserving perplexity, and integrates seamlessly with Flash Attention. Our analysis offers new insights into attention biases in parallel attention mechanisms and provides practical solutions to tackle these challenges. To mitigate the attention sink issue, we propose an attention calibration strategy that reduces biases, ensuring more stable long-range attention. Additionally, we introduce a chunk eviction strategy to efficiently manage ultra-long contexts on a single A100 80GB GPU. To further enhance efficiency, we propose a parallel KV cache eviction technique, which improves chunk throughput by 1.76×, thereby achieving a 23.50× acceleration in the prefilling stage with negligible performance loss due to attention calibration. Furthermore, ParallelComp achieves 91.17% of GPT-4’s performance on long-context tasks using an 8B model trained on 8K-length context, outperforming powerful closed-source models such as Claude-2 and Kimi-Chat.
@@ 1 Introduction
Processing long contexts is a fundamental capability of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024). Rotary position embeddings (RoPEs) (Su, 2023) are adopted to train LLMs due to their ability to encode relative positions. However, achieving length extrapolation through retraining or fine-tuning the entire model remains a challenging problem (Press et al., 2022; Chi et al., 2022; Peng et al., 2023; Fu et al., 2024), especially in resource-constrained situations.
Figure 1: Upper: The distribution of two types of attention. Lower: The heatmaps of two types of attention.
To address this, many researchers explored training-free approaches for context length extrapolation. One prominent direction involves modifying positional encodings to facilitate length extrapolation (LocalLLaMA, 2023b, a; An et al., 2024). Another promising strategy employs text chunking techniques (An et al., 2024; Xiao et al., 2024; Ratner et al., 2022; Zhu et al., 2024), which form a type of parallel attention mechanism. These methods often build upon the classic RoPE mechanism to enhance the model’s length extrapolation capability. However, both approaches suffer from the attention sink phenomenon (Liu et al., 2024; Xiao et al., 2023; Han et al., 2024; Gu et al., 2024), where high attention scores tend to be assigned to the first few tokens or the last few tokens in the input sequence, making the model fail to capture essential contextual information. Figure 1 illustrates this phenomenon and contrasts the attention distributions in normal and parallel attention. In the upper part of Figure 1, we observe the distribution of attention scores across tokens. In the normal attention distribution (left), the attention is heavily skewed towards the first and last few tokens, a common issue that limits the model’s ability to capture long-range dependencies. In contrast, the parallel attention distribution (right) attempts to spread the attention more evenly, but it still shows noticeable patterns of concentration in certain regions, indicating room for improvement in handling long-contexts effectively. The corresponding heatmaps (c) and (d) further reveal the attention matrices for both normal and parallel attention, where (c) shows the typical concentration of attention in the initial tokens, and (d) illustrates the fragmentation and partial concentration in parallel attention. In this paper, we focus on the unique phenomenon of attention sink in parallel attention for length extrapolation and explore solution to tackle this challenge.
Apart from the attention sink phenomenon, we also observe some unique biases in parallel attention. The impact of these attention biases on the model’s length extrapolation ability is still under exploration. To promote the understanding of the effects of special attention biases on parallel attention, we propose the following questions in this paper: Q1: What types of attention patterns can be summarized? Q2: Is there any difference between the attention bias in parallel attention and the attention bias in classical attention? Q3: Can the calibration strategy alleviate attention bias?
@@
To address the above questions, we propose ParallelComp, a parallel long-context compression method to extrapolate length. While maintaining high throughput, we extrapolate the length from 4k to 128k on a single GPU, without increasing perplexity. Overall, our contributions are as follows:
@@
  • We propose ParallelComp, a novel training-free method that enables efficient length extrapolation for LLMs. Our approach extrapolates from a 4K context length to up to 128K tokens on a A100 80GB GPU.
  • We conduct a detailed analysis of the unique attention biases in parallel attention mechanisms and introduce an attention calibration strategy that effectively alleviates these biases, recovering the performance loss.
@@
  • Our chunk eviction strategy enables parallel attention to handle contexts exceeding 128K on a single A100, while parallel KV cache eviction further enhances chunk throughput by 1.76x, thereby achieving a 23.50x acceleration in the prefilling stage with negligible performance loss due to attention calibration.
@@
  • Experiments show that ParallelComp achieves 91.17% of GPT-4’s performance on ultra-long context tasks with an 8B model trained on 8K-length context, surpassing powerful closed-source models such as Claude-2 and Kimi-Chat.
@@
Figure 2: Overview of ParallelComp. Parallel Attention – The input sequence is split into multiple chunks based on the model’s maximum context length. Each chunk undergoes local attention computation independently, and the self-information score of the query is calculated. Parallel KV Cache Eviction – Based on the self-information score, low-score tokens (marked in yellow, $R_{l}$) and attention bias tokens (marked in red, $R_{h}$) are selectively evicted to optimize memory usage and attention bias. Global Attention – The remaining KV caches are ranked by self-information, and less relevant chunks are discarded. The selected chunks are then concatenated, and a global attention operation is applied to ensure comprehensive information aggregation before the final autoregressive decoding stage.
@@ 2 Related Work
Positional Encoding.
Existing absolute positional encoding (APE) (Vaswani, 2017; Devlin, 2018) incorporates either fixed or learnable positional encodings into input representations through vector addition. However, APE faces challenges when dealing with long contexts. To overcome these limitations, relative positional encoding (RPE) methods—such as rotary and additive positional encodings (Su, 2023; Su et al., 2024; Press et al., 2021)—are developed to offer improved generalization in longer contexts. In addition, some data-dependent positional encoding methods (Golovneva et al., 2024; Zheng et al., 2024) are gaining widespread attention. These positional encodings are generated based on the input data.
Some reecent works (Press et al., 2022; Chi et al., 2022; Li et al., 2023) also focus on designing better positional encodings to enhance the pre-trained model’s capability for length extrapolation. Another line of works (Peng et al., 2023; bloc97, 2023; Chen et al., 2023a) focus on enhancing the LLM’s length extrapolation capability by fine-tuning.
Furthermore, there are two categories of training-free extrapolation methods. The first category, such as LocalLLaMA (2023b); Chen et al. (2023b); LocalLLaMA (2023a); Chen et al. (2024), directly modifies positional encodings to enable extrapolation or interpolation, aiming to enhance the model’s length extrapolation capability. The second category (An et al., 2024; Xiao et al., 2024; Ratner et al., 2022; Zhu et al., 2024) achieves extrapolation solely by reusing positional encodings. In this work, we focus primarily on the training-free setting.
@@
A series of studies (Gu et al., 2024; Xiao et al., 2023; Liu et al., 2024) reveal the phenomenon of attention sink, where certain tokens in the sequence (referred to as sink tokens) consistently receive abnormally high attention scores. When the sequence length increases, the attention scores in the middle of the sequence are significantly lower compared to the beginning and the end. This often prevents the model from focusing on the correct parts of long sequences. Xiao et al. (2023) attributes this phenomenon to the softmax function, which forces the query to assign attention scores to all preceding tokens, even when the preceding tokens lack essential information, resulting in high scores being assigned to initial tokens. Gu et al. (2024) proposes a simple solution by replacing the softmax attention with alternative attention mechanisms (e.g., unnormalized sigmoid attention) to alleviate this dependency. Chen et al. (2024) alleviates this phenomenon by simply removing certain low-frequency components from RoPE. Yu et al. (2024) deals with this issue by calibrating the attention distribution. In our work, we focus primarily on three phenomena of attention bias within parallel attention: the attention sink at the beginning of the input, the attention sink at the end of the input (i.e. recency bias Peysakhovich & Lerer 2023), and the scattered attention in the middle of the input (Yu et al., 2024). These biases provide insights into how LLMs utilize parallel contextual information.
@@ 3 Method
In this section, we introduce ParallelComp, our proposed approach for achieving efficient long-context inference. Then we introduce its unique bias phenomenon. Figure 2 offers a high-level overview of ParallelComp.
@@ 3.1 ParallelComp
Inspired by previous studies (Chen et al., 2023c; An et al., 2024), we split the text into chunks according to the model’s maximum context size and concatenate them with the input query for parallel encoding. This step is typically performed using local attention. For a given input sequence $X\in\mathbb{R}^{N\times d}$, the sequence is divided into $C=\lceil N/w\rceil$ chunks, each containing at most $w$ tokens (the maximum context length of each chunk). Let $f_{Q}(\cdot)$, $f_{K}(\cdot)$, and $f_{V}(\cdot)$ represent the linear transformation functions for query, key, and value projections, respectively. Then, the attention computation is performed parallelly within each chunk [^1] We use multi-head attention in implementation but omit the head information for simplicity in our description.:
@@
To ensure that the computation of parallel attention can be performed on a single 80GB A100 GPU, we design a chunk eviction strategy to control memory overhead as shown in Figure 2 step 3. Inspired by Ye et al. (2023), we introduce a chunk eviction mechanism that leverages the self-information of the query tokens $X^{q}$ to further enhance parallel processing efficiency. This mechanism utilizes an online priority queue to manage memory, retaining only the most relevant chunks, thereby enhancing language modeling. For a given chunk $c$, the self-information score for the query tokens $X^{q}$ is calculated as follows:
@@
$S=\{c\mid I_{c}\leq\tau\},$ (4)
where $\tau$ is a threshold that determines whether a chunk will be selected or not. The selected set of chunks is stored in a fixed-size priority queue to ensure that the prefilling stage remains within the memory limit.
@@
To further increase the throughput of chunks, we design a KV cache eviction strategy as shown in Figure 2 step 2. We utilize Flash Attention to perform fast attention computation. Since it is hard to obtain the complete attention distribution from Flash Attention (Dao, 2023), we use the local attention of $X^{q}$ to quickly estimate tokens with relatively low attention scores and evict them in advance before sending them into Flash Attention:
$S_{c,j}=\sum_{i=1}^{w_{q}}A^{c}_{\mathfrak{l}}(X^{q}_{i},X^{c}_{j}),\quad j=1,2,...,w,$ (5)
where $S_{c}\in\mathbb{R}^{w}$ represents the cumulative attention score for each token in the chunk and $A^{c}_{\mathfrak{l}}$ is the local attention score between the $i$-th token of the query $X^{q}$ and the $j$-th token of the chunk $X^{c}$, $w_{q}$ denotes the length of the query sequence. The cumulative attention score aggregates the attention distributions from each token in the query to each token in the chunk, thereby measuring the relevance of each token in the chunk to the query $X^{q}$. Tokens with low cumulative attention scores within the chunk are evicted, and the retained tokens are used to form the compressed KV cache.
$K_{r}=K_{x}[R_{l}],\quad V_{r}=V_{x}[R_{l}],$ (6)
where $K_{x}$ and $V_{x}$ represent the KV cache of the input chunk, and $R_{l}$ denotes the set of indices corresponding to the evicted tokens with low attention scores. The notation []delimited-[][$\cdot$][ ⋅ ] indicates indexing into $K_{x}$ and $V_{x}$ to evict only the tokens corresponding to the indices in $R_{l}$. $K_{r}$ and $V_{r}$ denote the retained KV cache. The compression strategy of the KV cache typically helps reduce memory overhead while increasing chunk throughput, but it often exacerbates attention bias. Next, we introduce a simple and effective strategy to calibrate attention distribution.
@@
To alleviate the attention bias exacerbated by Parallel KV Cache Eviction, we design another token eviction strategy using Eq. 5. Specifically, we evict tokens with excessively high attention scores. Let $R_{h}$ correspond to those with extremely high attention scores exceeding $\lambda$ (a manually-set threshold), then we have:
$K^{\prime}_{r}=K_{r}[R_{h}],\quad V^{\prime}_{r}=V_{r}[R_{h}].$ (7)
Evicting tokens with exceptionally high scores guarantees that the segmented computation of softmax in Flash Attention can produce calibrated attention distributions. We will thoroughly investigate the impact of this calibration method on the attention distribution in Section 4.
@@
After obtaining the attention outputs for each chunk, we concatenate the key-value (KV) caches from all chunks into a unified representation. Specifically, the concatenated KV cache is given by:
@@
where $K^{X^{c}}$ and $V^{X^{c}}$ are the key and value projections of the $c$-th chunk.
@@
where $F_{\mathfrak{g}}\in\mathbb{R}^{w_{q}\times d}$ represents the globally updated features for the query chunk. Finally, the updated global representation is passed through the feedforward and autoregressive decoding stages, enabling the model to generate outputs while leveraging information from all chunks efficiently.

### v2

@@ Abstract
Extrapolating ultra-long contexts (text length $>128$K) remains a major challenge for large language models (LLMs), as most training-free extrapolation methods are not only severely limited by memory bottlenecks, but also suffer from the attention sink, which restricts their scalability and effectiveness in practice. In this work, we propose ParallelComp, a parallel long-context compression method that effectively overcomes the memory bottleneck, enabling 8B-parameter LLMs to extrapolate from 8K to 128K tokens on a single A100 80GB GPU in a training-free setting. ParallelComp splits the input into chunks, dynamically evicting redundant chunks and irrelevant tokens, supported by a parallel KV cache eviction mechanism. Importantly, we present a systematic theoretical and empirical analysis of attention biases in parallel attention—including the attention sink, recency bias, and middle bias—and reveal that these biases exhibit distinctive patterns under ultra-long context settings. We further design a KV cache eviction technique to mitigate this phenomenon. Experimental results show that ParallelComp enables an 8B model (trained on 8K context) to achieve $91.17\%$ of GPT-4’s performance under ultra-long contexts, outperforming closed-source models such as Claude-2 and Kimi-Chat. We achieve a 1.76x improvement in chunk throughput, thereby achieving a 23.50x acceleration in the prefill stage with negligible performance loss and pave the way for scalable and robust ultra-long contexts extrapolation in LLMs. We release the code at https://github.com/menik1126/ParallelComp.
@@ 1 Introduction
Extrapolating long-contexts is a core capability of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023; Dubey et al., 2024). Achieving this requires effective mechanisms for modeling positional relationships over extended sequences. Rotary position embedding (RoPE) (Su, 2023) are commonly employed in LLMs due to their ability to efficiently encode relative positional information. However, extrapolating to lengths beyond the training range during inference remains a significant challenge. While retraining or fine-tuning the entire model is one possible solution, it is not always feasible, particularly in resource-constrained environments (Press et al., 2022; Chi et al., 2022; Peng et al., 2023; Fu et al., 2024). This limitation prevents RoPE-based models from generalizing to longer sequences without additional training, making them less adaptable to scenarios requiring inference on inputs of unforeseen lengths.
Figure 1: Upper: Distribution of two types of attention. Lower: Heatmaps of the two types of attention. The attention head shown here represents the distribution change between normal attention and parallel attention for the 21st head in layer 1, using the Llama-2-7b-chat-hf model.
To address this, many researchers explore training-free approaches for context length extrapolation. One prominent direction involves incorporating NTK-based methods (LocalLLaMA, 2023b, a; bloc97, 2023), which distinguish dimensions of different frequencies by their wavelengths and apply tailored interpolation strategies to address the extrapolation challenge of position encoding. Another promising strategy leverages text chunking techniques (An et al., 2024; Xiao et al., 2024; Ratner et al., 2022; Zhu et al., 2024), which typically reuse position encodings across chunks to enable chunk-wise length extrapolation without retraining, thereby improving the model’s ability to generalize to longer contexts. However, both approaches suffer from the attention sink phenomenon (Liu et al., 2024; Xiao et al., 2023; Han et al., 2024; Gu et al., 2024), where high attention scores tend to be assigned to the first few tokens or the last few tokens in the input sequence. Figure 1 illustrates this phenomenon and compares the attention distributions between standard and parallel attention mechanisms. In the upper part of Figure 1, we observe the distribution of attention scores across tokens. In the standard attention distribution (left), corresponding to the heatmaps (c) below the image, attention is disproportionately focused on the first and last few tokens. In contrast, the parallel attention distribution (right), shown in the heatmaps (d) below the image, attempts to distribute attention more evenly. However, it still shows noticeable concentrations in certain regions and exhibits a multi-peak distribution, indicating that the nature of the attention sink differs fundamentally between the two mechanisms. The phenomenon within the length extrapolation mechanism in the parallel attention mechanism remains unexplored. This paper focuses on addressing the memory limitations encountered during length extrapolation and provides a detailed analysis of the unique attention sink that emerges in parallel attention. Specifically, we extrapolate the length by chunking the input and reusing position encodings, applying parallel compression of the KV cache of the chunks and tokens to resolve memory bounds, and exploring the unique phenomenon of the attention sink in parallel attention while attempting to adopt a token eviction strategy to mitigate this bias. To promote the understanding of the effects of special attention sink on parallel attention, we propose the following questions in this paper: Q1: What types of attention patterns can be summarized? Q2: Is there any difference between the attention bias in parallel attention and the attention bias in classical attention? Q3: Can the calibration strategy alleviate attention bias? Q4: Does our parallel compression strategy effectively support length extrapolation?
@@
To address the above questions, we propose ParallelComp, a parallel long-context compression method to extrapolate length. While maintaining high throughput, we extrapolate the length from 8K to 128K on a single GPU, with almost no performance loss. Overall, our contributions are as follows:
@@
  • We propose ParallelComp, a novel training-free method that enables efficient length extrapolation for LLMs — scaling from 8K to up to 128K tokens on a single A100 80GB GPU — by addressing attention biases through an effective attention calibration strategy.
@@
  • To overcome limitations in parallel attention, we introduce a chunk eviction mechanism and parallel KV cache eviction, allowing processing of contexts beyond 128K tokens and achieving a 1.76x throughput improvement and a 23.50x speedup in the prefilling stage, with negligible performance loss.
@@
  • Experiments demonstrate that ParallelComp achieves 91.17% of GPT-4’s performance on ultra-long context tasks using an 8B model trained on only 8K-length context, outperforming strong proprietary models such as Claude-2 and Kimi-Chat.
@@
Figure 2: Overview of ParallelComp. Parallel Attention – The input sequence is split into multiple chunks based on the model’s maximum context length. Each chunk undergoes local attention computation independently, and the self-information score of the query is calculated. Parallel KV Cache Eviction – Based on the self-information score, low-score tokens (marked in yellow, $R^{h}_{L}$) and attention bias tokens (marked in red, $R^{h}_{H}$) are selectively evicted to optimize memory usage and attention bias. Global Attention – The remaining KV caches are ranked by self-information, and less relevant chunks are discarded. The selected chunks are then concatenated, and a global attention operation is applied to ensure comprehensive information aggregation before the final autoregressive decoding stage.
@@ 2 Related Work
Position encoding.
Existing absolute position encoding (APE) (Vaswani, 2017; Devlin, 2018) incorporates either fixed or learnable position encodings into input representations through vector addition. However, APE faces challenges when dealing with long-contexts. To overcome these limitations, relative position encoding (RPE) methods—such as rotary and additive position encodings (Su, 2023; Su et al., 2024; Press et al., 2021)—are developed to offer improved generalization in longer contexts. In addition, some data-dependent position encoding methods (Golovneva et al., 2024; Zheng et al., 2024) are gaining widespread attention. These position encodings are generated based on the input data.
Some works (Press et al., 2022; Chi et al., 2022; Li et al., 2023a) also focus on designing better position encodings to enhance the pre-trained model’s capability for length extrapolation. Another line of works (Peng et al., 2023; bloc97, 2023; Chen et al., 2023a) focus on enhancing the LLM’s length extrapolation capability by fine-tuning.
Furthermore, there are two categories of training-free extrapolation methods. The first category, such as LocalLLaMA (2023b); Chen et al. (2023b); LocalLLaMA (2023a); Chen et al. (2024), directly modifies position encodings to enable extrapolation or interpolation, aiming to enhance the model’s length extrapolation capability. The second category (An et al., 2024; Xiao et al., 2024; Ratner et al., 2022; Zhu et al., 2024) achieves extrapolation solely by reusing position encodings. In this work, we focus primarily on the training-free setting.
@@
A series of studies (Gu et al., 2024; Xiao et al., 2023; Liu et al., 2024) reveal the phenomenon of attention sink, where certain tokens in the sequence (referred to as sink tokens) consistently receive abnormally high attention scores.
When the sequence length increases, the attention scores in the middle of the sequence are significantly lower compared to the beginning and the end. Some work (Xiao et al., 2023; Zhang et al., 2023; Wan et al., 2024; Xiong et al., 2024) find that the tokens of attention sink can effectively aggregate information and use it to efficiently compress the KV cache. But in ultra-long contexts, this may prevent the model from focusing on the correct parts of long sequences. Xiao et al. (2023) attributes this phenomenon to the softmax function, which forces the query to assign attention scores to all preceding tokens, even when the preceding tokens lack essential information, resulting in high scores being assigned to initial tokens. Gu et al. (2024) proposes a simple solution by replacing the softmax attention with alternative attention mechanisms (e.g., unnormalized sigmoid attention) to alleviate this dependency. Chen et al. (2024) alleviates this phenomenon by simply removing certain low-frequency components from RoPE. Yu et al. (2024) deals with this issue by calibrating the attention distribution. In our work, we focus primarily on three phenomena of attention bias within parallel attention: the attention sink at the beginning of the input, the attention sink at the end of the input (i.e. recency bias Peysakhovich & Lerer 2023), and the scattered attention in the middle of the input (Yu et al., 2024). These biases provide insights into how LLMs utilize parallel contextual information.
@@ 3 Method
In this section, we introduce ParallelComp, our proposed approach for achieving efficient long-context compression during extrapolation. We then discuss its unique bias phenomenon. Figure 2 offers a high-level overview of ParallelComp.
@@ 3.1 ParallelComp
Inspired by previous studies (An et al., 2024), we divide the text into chunks based on the model’s maximum context size and concatenate them with the input query for parallel encoding. This step is typically carried out using local attention. For a given input sequence $X\in\mathbb{R}^{N\times d}$, the sequence is split into $C=\lceil N/w\rceil$ chunks, each containing at most $w$ tokens (the maximum context length per chunk). Let $f_{Q}(\cdot)$, $f_{K}(\cdot)$, and $f_{V}(\cdot)$ denote the linear transformation functions for the query, key, and value projections, respectively. Attention computation is then performed within each chunk:
@@
To ensure that the computation of parallel attention can be performed on a single 80GB A100 GPU, we design a chunk eviction strategy to control memory overhead as shown in Figure 2 step 3. Inspired by Li et al. (2023b); Jiang et al. (2023), we introduce a chunk eviction mechanism that leverages the self-information of the query tokens $X^{q}$ to further enhance parallel processing efficiency. This mechanism employs an online priority queue to manage memory, retaining only the most relevant chunks with the lowest perplexity, thereby improving language modeling. For a given chunk $c$, the self-information score for the query tokens $X^{q}$ is calculated as follows:
@@
$S=\{c\mid c\leq\epsilon\},$ (4)
where $\epsilon$ is a threshold that determines whether a chunk will be selected or not. The selected chunks are sorted based on $I_{c}$, and only a fixed number of top-ranked chunks are retained. These selected chunks are stored in a fixed-size priority queue to ensure that the prefilling stage remains within the memory limit.
@@
To further improve chunk throughput, we propose a KV cache eviction strategy, as illustrated in Step 2 of Figure 2. We leverage Flash Attention (Dao, 2023) for efficient attention computation. Prior to performing local attention, we use the cumulative scores of $X^{q}$ to quickly identify tokens with relatively low attention importance and evict them. Specifically, the local attention score $A^{c}_{\mathfrak{l}}$ between the $i$-th token of the query $X^{q}_{i}$ and the $j$-th token of the chunk $X^{c}_{j}$ is calculated as:
$A^{c}_{\mathfrak{l(i,j)}}=f(X^{q}_{i},X^{c}_{j}),$ (5)
where $f$ represents the matrix multiplication used to compute the attention score between $X^{q}_{i}$ and $X^{c}_{j}$. Then, the cumulative attention score for the $j$-th token in the $c$-th chunk is computed by summing the local attention scores over all tokens in the query chunk:
$S_{c,j}=\sum_{i=1}^{w_{q}}A^{c}_{\mathfrak{l(i,j)}},\quad j=1,2,...,w,$ (6)
where $S_{c,j}\in\mathbb{R}^{w}$ denotes the cumulative attention score for the $j$-th token in the $c$-th chunk, and $w_{q}$ is the length of the query chunk $X^{q}$. The cumulative attention score aggregates the attention distributions from each token in the query to each token in the chunk, thereby measuring the relevance of each token in the chunk to the query $X^{q}$. Tokens with low cumulative attention scores within the chunk are evicted, and the retained tokens are used to form the compressed KV cache:
$K^{h}_{r}=K^{h}_{x}[{R^{h}_{L}}],\quad V^{h}_{r}=V^{h}_{x}[R^{h}_{L}],$ (7)
where $K^{h}_{x}$ and $V^{h}_{x}$ represent the KV cache of the $h$-th head of the input chunk, and $R^{h}_{L}$ denotes the set of indices corresponding to the evicted tokens in the $h$-th head with low attention scores. The notation []delimited-[][$\cdot$][ ⋅ ] indicates indexing into $K^{h}_{x}$ and $V^{h}_{x}$ to evict only the tokens corresponding to the indices in $R^{h}_{L}$. $K^{h}_{r}$ and $V^{h}_{r}$ denote the retained KV cache. The compression strategy of the KV cache typically helps reduce memory overhead while increasing chunk throughput, but it often exacerbates attention bias. Next, we introduce a simple and effective strategy to calibrate attention distribution.
@@
To mitigate the attention bias exacerbated by parallel KV cache eviction, we propose an alternative token eviction strategy based on Eq. 6. Specifically, we evict tokens with excessively high attention scores. Let $R^{h}_{H}$ represent the tokens with attention scores of $h$-th head exceeding a manually-set threshold $\lambda$. Thus, we have:
$K^{h}_{r^{\prime}}=K^{h}_{x}[{R^{h}_{H}}],\quad V^{h}_{r^{\prime}}=V^{h}_{x}[R^{h}_{H}].$ (8)
Evicting tokens with exceptionally high scores guarantees that the attention mechnism can produce calibrated attention distributions. We will thoroughly investigate the impact of this calibration method on the attention bias in Section 4.
@@
After obtaining the attention outputs for each chunk, we concatenate the KV caches from all chunks into a unified representation. Specifically, the concatenated KV cache is given by:
@@
where $K^{X^{c}}$ and $V^{X^{c}}$ are the KV caches of the $c$-th chunk.
@@
where $F_{\mathfrak{g}}\in\mathbb{R}^{w_{q}\times d}$ represents the globally updated features for the query chunk. Finally, the updated global representation is passed through the decoding stages, enabling the model to generate outputs while effectively leveraging information from all chunks.